(6.1) Find the values of a, b and c such the matrix below is skew symmetric.

0 0 d

0 2a − 3b + c 3a − 5b + 5c

2 0 5a − 8b + 6c


(6.2) Give an example of a skew symmetric matrix.

(6.3) Prove that A (4) 2 is symmetric whenever A is skewsymmetric.

(6.4) Determine an expression for det(A) in terms of det(A (T ) if A is a square skewsymmetric.

(6.5) Assume that A has an odd number of rows and also an odd number of columns. In this particular case, show that det(·) is an odd function


1
Expert's answer
2021-05-25T15:41:06-0400

1)

Given matrix is skew symmetric matrix:

"A'-A"

"\begin{pmatrix}
 0 & 0 &2\\
 0 & 2a-3b+c &0\\
d & 3a-5b+5c &5a-8b+6c\\
\end{pmatrix}=-\begin{pmatrix}
 0 & 0 &d\\
 0 & 2a-3b+c &3a-5b+5c\\
2 & 0 &5a-8b+6c\\
\end{pmatrix}"

On comparing we get:

"d=2"

"2a-3b+c=-2a+3b-c"

"\implies 2a-3b+c=0"

"\implies 3a-5b+5c=0"

"\implies -5a+8b-6c=0"

Solving above equations we have:

"a=0,b=0,c=0"


2)

Skew symmetric matrix is

"\begin{pmatrix}
 0 & 1&-2 \\
 -1 & 0&3 \\
 2&-3&0
\end{pmatrix}"


3)

Let "B=A^2"

if "A=A^T\implies AA^T=A^2=B"

"B^T=(A^TA)^T=A^T(A^T)^T=A^TA=A^2=B"

"B" is symmetric.

Therefore "A^2" is symmetric.


4)

"det(A^T)=|A^T|"

"|-A|=(-1)^n|A|=-|A|"

"det(A^T)=-det(A)"


5)

Given that A matrix is a square matrix,

 let n=3 be the order

"|-A|=(-1)^3|A|=-|A|"

"|-A|=-|A|"

So, "det()" is an odd function.